If the price of a slice of pizza rises from $2.50 to $3, and quantity demanded falls from 10,000 slices to 7,400 slices, using the formula for arc price elasticity

a)         What is the level of elasticity?

b)         What is the interpretation the result


1
Expert's answer
2021-07-22T10:00:41-0400

(a) By the definition of the arc price elasticity, we have:


"E_d=\dfrac{\%\Delta Q}{\%\Delta P}=\dfrac{\dfrac{Q_2-Q_1}{0.5(Q_2+Q_1)}}{\dfrac{P_2-P_1}{0.5(P_2+P_1)}},""E_d=\dfrac{\dfrac{7400-10000}{0.5\cdot(7400+10000)}}{\dfrac{\$3-\$2.50}{0.5\cdot(\$3+\$2.50)}}=-1.64."


(b) As we can see from calculations, "E_d>1". Therefore, the demand is elastic.